OpenStudy (anonymous):

6 divided by 5square root 3

OpenStudy (anonymous):

\[6/\sqrt{3}\] Just you a calculator and you get 3.46

OpenStudy (radar):

\[6/(5\sqrt{3})\] is this your problem?

OpenStudy (radar):

If so you should get\[(4/5)\times \sqrt{3}=.8\sqrt{3}=1.38564\]